Last year Ann Arbor Corp had $250,000 of assets (which equals total invested capital), $305,000 of sales, $20,000 of net income, and a debt-to-total-capital ratio of 37.5%. The new CFO believes that a new computer program will enable the company to reduce costs and thus raise net income to $33,000. The firm finances using only debt and common equity. Assets, total invested capital, sales, and the debt to capital ratio would not be affected. By how much would the cost reduction improve the ROE? Do not round your intermediate calculations.

a.

6.74%

b.

8.82%

c.

8.32%

d.

8.15%

e.

8.57%
Answer #1

Ans)

Assets= Total invested capital = $250,000

Debt Rate = 37.5%

Debt= capital X Debt % = 93750

Equity= Assets – Debt = 250,000 - 93750 = 156250

Sales = $305000

Old NI = $20000

New NI =$33000

New ROE= New NI/Equity =33000/156250 = 21.12%

Old ROE= Old NI/Equity= 20,0000/156250 = 12.8%

Increase in ROE= New ROE- Old ROE = 21.12 - 12.8 = 8.32%